# Hilbert space methods for approximating multi-output latent variable Gaussian processes

### v2

@@ Acknowledgments
Figure 20: Periodic data with lower oscillations scenario: posterior bias and SD on recovery of latent inputs for all fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
Figure 21: Periodic data with higher oscillations scenario: posterior bias and SD on recovery of latent inputs for all fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
F: Hyperparamter estimation (additional figures)
@@
Figure 24: Periodic data with lower oscillations scenario: RMSE on recovery of GP hyperparameters for exact GP and HSGP fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
Figure 25: Periodic data with higher oscillations scenario: RMSE on recovery of GP hyperparameters for exact GP and HSGP fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
G: List of genes for the case study
In our real-world case study for cell cycle data, we use the data from all the cells along with a selection of 12 influential genes (see supplementary materials Section G for the list) based on the recommendations of Mahdessian et al., (2021). The list of genes used are: CCNA2, CCNB1, BIRC5, TOP2A, PLK1, NDC80, CDCA8, MKI67, CDC6, CENPF, KRT17, RRM2. These recommendations were primarily concerned with high amount of RNA activities denoted by their counts.

### v3

@@ Acknowledgments
Figure 27: Periodic data scenario (lower oscillations): posterior bias and SD on recovery of latent inputs for all fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
Figure 28: Periodic data scenario (higher oscillations): posterior bias and SD on recovery of latent inputs for all fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
G: Checking robustness of HSGPs with wide priors (additional figures)
Figure 29: Matern 3/2 scenario (special case): posterior bias and SD on recovery of latent inputs for all fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions. VIGP (default) shows the default Pyro implementation and VIGP (prior) has fixed GP hyperparameters informed by the true sampling distributions.
Figure 30: Matern 5/2 scenario (special case): posterior bias and SD on recovery of latent inputs for all fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions. VIGP (default) shows the default Pyro implementation and VIGP (prior) has fixed GP hyperparameters informed by the true sampling distributions.
H: Hyperparameter estimation (additional figures)
The GP hyperparameter estimation results for the Matern 3/2, 5/2 and the periodic data scenarios (both lower and higher oscillations) are presented here. The results are qualitatively similar across all the simulation scenarios.
@@
Figure 33: Periodic data scenario (lower oscillations): RMSE on recovery of GP hyperparameters for exact GP and HSGP fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
Figure 34: Periodic data scenario (higher oscillations): RMSE on recovery of GP hyperparameters for exact GP and HSGP fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
I: List of genes for the case study
In our real-world case study for cell cycle data, we use the data from all the cells along with a selection of 12 influential genes based on the recommendations of Mahdessian et al. (2021). The list of genes used are: CCNA2, CCNB1, BIRC5, TOP2A, PLK1, NDC80, CDCA8, MKI67, CDC6, CENPF, KRT17, RRM2. These recommendations were primarily concerned with high amount of RNA activities denoted by their counts.